OpenStudy (anonymous):

Write and solve an equation to find the value of the variable. 5.2, 8.3, 8.5, 7.7, 7.8, y; mean = 7.1 5.1 5.3 5.8 3.9

OpenStudy (xapproachesinfinity):

the mean is the average the average is the sum of all number divided by how many number there are

OpenStudy (anonymous):

ok thanks

OpenStudy (xapproachesinfinity):

so you do 5.2+8.3+....+y/6=7.1

OpenStudy (xapproachesinfinity):

Not only y/6 all the sum divided by 6 ok

OpenStudy (anonymous):

5.2 + 8.5

OpenStudy (xapproachesinfinity):

what is that?

OpenStudy (xapproachesinfinity):

i said sum everything up

OpenStudy (anonymous):

ok mybad

OpenStudy (xapproachesinfinity):

ok work-_^

OpenStudy (anonymous):

i got 7.34

OpenStudy (xapproachesinfinity):

for the sum?

OpenStudy (xapproachesinfinity):

(5.2+ 8.3+8.5+7.7+7.8+y)/6 = 7.1

OpenStudy (xapproachesinfinity):

is this what you did

OpenStudy (anonymous):

the answer is A. 5.1. The way I solved it was to work backwards. First I multiplied 7.1 by 6 to get the total amount of all the numbers added together. Next I subtracted the total of the numbers we already know (e.g. 5.2, 8.3, 8.5, 7.7, 7.8). The remaining amount is the answer. Let me know if you don't understand what I said so I can explain further.

OpenStudy (anonymous):

Also the exact equation is 7.1 X 6 - (5.2 + 8.3 + 8.5 + 7.7 + 7.8) = x